OpenStudy (anonymous):

consider the piecewise h(x) = -2/3(x+1) x<-1 -4x-4 -1≤ x≤ 0 x-8 1

OpenStudy (aum):

Use -2/3(x +1) IF x < -1 Use -4x - 4 IF -1≤ x≤ 0 Use x - 8 IF 1 < x < 4 h(-3) = ? x = -3. Which one of the above should you use?

OpenStudy (anonymous):

okay so h(-3) = 4/3

OpenStudy (aum):

Find h(-3) When x = -3, it is less than -1. So you use the first one: Use -2/3(x +1) IF x < -1 h(x) = -2/3 * (x + 1) put x = -3 h(-3) = -2/3 * (-3 + 1) = -2/3 * (-2) = 4/3 Next find h(0). That is, what is h when x = 0. Look in my very first reply and select which one to use when x = 0.

OpenStudy (anonymous):

okay so h(0) is -4 and h(3) is -5

OpenStudy (anonymous):

so what is the domain of h(x)?

OpenStudy (aum):

Correct on h(0) and h(3).